OpenStudy (anonymous):

Try this one: There are \(100\) transistors in a box of which some are defective. At random, two transistors are consecutively taken out without replacement. A scientist wants to know the chances of exactly one of them being good and the other being defective. What number of defective transistors takes this chance below \(30 \% \)?

OpenStudy (sasogeek):

15... :/

OpenStudy (anonymous):

Not that hard, give it a quick try guys!!

OpenStudy (anonymous):

I am sorry saso, that is not the correct answer.

OpenStudy (sasogeek):

is the answer above that number or below? more or less...

OpenStudy (anonymous):

above, more than the double of that number actually.

OpenStudy (sasogeek):

35

OpenStudy (anonymous):

Is the chance of it being defective 30%?

OpenStudy (anonymous):

no, only below.

OpenStudy (sasogeek):

hmmm seems harder than i thought

OpenStudy (sasogeek):

i don't know how to solve this but i'm trying to use logic to find my way out... i need a couple more hints to know if i'm close to the answer... you mind helping me out?

OpenStudy (anonymous):

sure I would love to, where are you stuck ?

OpenStudy (sasogeek):

i copied what i was typing to the clipboard, i'll paste it later, i'm being called. brb

OpenStudy (anonymous):

sure.

OpenStudy (dumbcow):

Let d be the number of defective transistors. probability of being defective is d/n, where n is remaining transistors probability of being good is (100-d)/n, where n is remaining transistors There are 2 possibilities, either the 1st transistor is good or it is defective and the 2nd is the opposite \[P = \frac{d}{100}*\frac{100-d}{99} + \frac{100-d}{100}*\frac{d}{99}\] \[P = \frac{2d(100-d)}{9900}\] set P = 0.3 and solve for d \[0.3 > \frac{2d(100-d)}{9900}\] \[1485 > 100d-d^{2}\] \[d^{2} -100d +1485 > 0\] \[d < 18.14\] \[d>81.86\]

OpenStudy (sasogeek):

i think the above statements make sense, the early parts falls within the logic i was trying to construct but the figures are different... oh well :)